# Quantum synchronization between two strongly driven YIG spheres mediated via a microwave cavity

### v1

@@ I INTRODUCTION
While extensively studied in classical contexts, recent research has extended this field into the quantum domain buvca2022algebraic ; e26050415 . Synchronization has been observed or predicted in a variety of quantum systems, ranging from subatomic particles to mechanical resonators manzano2013avoiding ; manzano2013synchronization and electrodynamic setups buvca2022algebraic ; eneriz2019degree . These systems often involve multiple interacting modes governed by nonlinear dynamics and subject to strong external drives. For instance, recent experiments demonstrated quantum synchronization of a single trapped ion qubit with an external signal zhang2023quantum . Quantum synchronization has been investigated also across various other platforms, namely, van der Pol oscillators giorgi2012quantum , Josephson junction arrays wiesenfeld1996synchronization , spin torque nano-oscillators kaka2005mutual and optomechanical systems holmes2012synchronization ; zhang2012synchronization ; heinrich2011collective ; PhysRevLett.111.103605 ; shim2007synchronized . It has been shown that quantum fluctuations, nonclassical correlations, and phase transitions play critical roles in quantum synchronization phenomena buvca2022algebraic ; eneriz2019degree ; PhysRevA.102.032414 . Recently we have developed a theoretical framework to analyze quantum synchronization between two finite-size spin chains that interact indirectly through a central spin chain ghildiyal2025quantum . This configuration results in effective nonlinear coupling, with dynamics influenced by both linear and nonlinear dissipation.
@@
In this paper, we focus on quantum synchronization between two magnon modes. Magnons are quanta of spin-wave excitations arising from the collective precession of localized spins princep2017full ; PhysRevResearch.2.022027 ; ghasemian2023dissipative ; kusminskiy2021cavity . It is recently shown that propagating spin waves can mediate phase-tunable synchronization between spin Hall nano-oscillators kumar2025spin .
@@
Ability to couple the magnons with the cavity modes and qubits has further led to a plethora of work related to quantum information processing in cavity-magnonic systems YUAN20221 . This development made it possible to explore quantum synchronization in such platform. For example, it is shown that magnon squeezing can significantly enhance entanglement between two indirectly coupled magnon modes in a magnomechanical system harraf2025quantum . Yang and his coworkers demonstrated asymmetric quantum synchronization between two magnon modes in a two-sublattice antiferromagnet coupled to a cavity yang2022asymmetric , highlighting a strong dependence on cavity resonance and input current direction. We also note that the magnetostriction in a cavity magnomechanical system enables robust synchronization between two mechanical modes, via nonlinear dispersive coupling PhysRevResearch.5.043197 . Nonreciprocal quantum synchronization in antiferromagnet-cavity systems has been reported in yang2021implementationenhancementnonreciprocalquantum .
@@
The Yttrium Iron Garnet (YIG) spheres host well-defined magnon modes governed by magnetostatic solutions to the Landau-Lifshitz-Gilbert equation, thanks to its spherical structure. These spheres are often preferred in the studies of magnons due to its low magnetic damping and high coherence, enabling long-lived spin excitations. By placing the YIG spheres inside a high-quality microwave cavity, strong coupling between the magnon modes and the cavity photons can be achieved. This interaction leads to the formation of hybridized states known as cavity magnon polaritons, which are important for quantum technologies involving signal processing, coherent transduction and nonlinear dynamics zhang2017observation . In this paper, we focus on quantum synchronization between two magnon modes of two such spatially separated YIG spheres, placed inside a cavity. The cavity-mediated mechanism enables the emergence of quantum synchronization, even in the presence of dissipation and detuning. We analyze the effects of key parameters such as coupling strength, detuning and nonlinear interactions between the magnons, and identify the conditions for which a stable synchronization emerges. Unlike other relevant works tyagi2024noise ; bittner2025noise , we have employed the input-output formalism, instead of master equation technique, to describe the effect of decoherence.
@@ II MODEL AND EQUATION OF MOTION
We consider two macroscopic YIG spheres, trapped inside a single-mode microwave cavity. Each sphere hosts many magnon modes. The lowest-order mode, known as the Kittel mode, corresponds to the collective spin ground state, featuring uniform spin precession across the sample princep2017full . The higher-order modes exhibit non-uniform spatial distributions with specific angular and radial profiles. Here, we consider the YIG spheres in their respective Kittel mode. The cavity mode with the annihiliation operator $a$ interacts coherently with these magnon modes with respective annihilation operators $m_{1}$ and $m_{2}$. This setup allows us to access information about the magnon modes through the cavity.
Figure 1: Schematic representation of two YIG (Yttrium Iron Garnet) spheres, $\hat{m}_{1}$ and $\hat{m}_{2}$, coupled via a common microwave cavity mode, $\hat{a}$. The blue curved lines represent the cavity, while the magenta circles indicate the YIG spheres. Solid arrows depict the interaction between the cavity mode and the YIG spheres, while dashed arrows represent feedback interactions.
The cavity is driven by an electromagnetic field of frequency $\omega_{L}$. Typically, the magnetic component of this driving field couples with the magnon mode of the macroscopic ensemble inside the sphere, via Zeeman interaction. The interaction Hamiltonian between the magnon modes and the cavity can then be written as PhysRevB.94.224410 ; PhysRevLett.104.077202 .
@@
Here $\omega_{c}$ is the cavity resonance frequency and $\omega_{m_{i}}$ is the frequency of the $i$th magnon mode. The $\Omega_{c}$ and $\Omega_{i}$ denote the Rabi frequencies, describing the amplitudes of the driving fields with respective frequencies $\omega_{L}$ and $\omega_{d}^{(i)}$, respectively. The term containing the Kerr coefficient $K_{i}$ arises from the magnetocrystalline anisotropy in YIG, which introduces intensity-dependent frequency shifts proportional to the magnon population $\langle m_{i}^{\dagger}m_{i}\rangle$ PhysRevB.94.224410 ; ganthya2022bistability . This term leads to a nonlinear phase-inducing evolution PhysRevB.94.224410 ; PhysRevResearch.1.023021 of the magnons. The Hamiltonian (2) contains the terms for cavity self-energy and the cavity driving, while the Eq. (3) is a sum of magnon self-energy, nonlinear energy shift, and magnon driving term, respectively. In the rotating frame with respect to the driving frequencies $\omega_{L}$ and $\omega_{d}^{(i)}$, the explicit time-dependence in the Hamiltonian gets eliminated, leading to the following form of the total Hamiltonian:
@@
where $\Delta_{c}=\omega_{c}-\omega_{L}$ and , $\Delta_{i}=\omega_{m_{i}}-\omega_{d}^{(i)}$ are the detunings of the cavity mode and the $i$th magnon mode, respectively, from the corresponding driving field.
This Hamiltonian describes the dynamics of the system, capturing the interaction between the magnon modes and the cavity. Incorporating the interaction picture with respect to the energy terms containing $a^{\dagger}a$ and $m_{i}^{\dagger}m_{i}$ and using the Baker–Campbell–Hausdorff formalism, the modified Hamiltonian $H^{\prime}(t)$ can be expressed as $(\iota=\sqrt{-1})$
@@
$\quad-g_{i}m_{i}^{\dagger}A_{i}e^{-\iota t(B_{i}-C_{i})}m_{i}a+\Omega_{i}A_{i}e^{-\iota t(D_{i}+C_{i})}m_{i}^{2}\bigg{]},$
@@
where $A_{i}=2\iota K_{i}t$,  $i\in 1,2$.
@@
Figure 2: (a) Limit-cycle trajectories in the $\bar{q_{1}}\leftrightharpoons\bar{p_{1}}$ (red) and $\bar{q_{2}}\leftrightharpoons\bar{p_{2}}$ (blue) spaces, (b) Variation of the mean values $\bar{q_{1}}$ (red) and $\bar{q_{2}}$ (blue), (c) the mean values $\bar{p_{1}}$ (red) and $\bar{p_{2}}$ (blue). The parameters chosen are $\Omega_{1}=1$, $\Omega_{2}=1.00001$. (d) Limit-cycle trajectories in the $\bar{q_{1}}\leftrightharpoons\bar{p_{1}}$ (red) and $\bar{q_{2}}\leftrightharpoons\bar{p_{2}}$ (blue) spaces, (e) Variation of the mean values $\bar{q_{1}}$ (red) and $\bar{q_{2}}$ (blue), (f) the mean values $\bar{p_{1}}$ (red) and $\bar{p_{2}}$ (blue). The parameters chosen are $\Omega_{1}=1$, $\Omega_{2}=1.1$ and $\phi=0.1320$. All the other parameters are the same in all the above figures: $\Omega_{c}=1$, $\Delta_{1}=\Delta_{2}=0.001$, $\Delta_{c}=-0.2$, $g_{1}=g_{2}=0.1$, $K_{1}=K_{2}=10^{-10}$, $\gamma_{1}=\gamma_{2}=\gamma_{c}=0.1$. All the parameters are normalized with respect to $\Omega_{1}$.
Figure 3: (a) Variation of $S_{q}^{\phi}$, with respect to time $t$ and (b) variation of $\bar{S}_{q}^{\phi}$, with respect to the mean phonon number $\bar{n}_{m}$ of the environment. The other parameters are $g_{1}=g_{2}=0.1$, $K_{1}=K_{2}=10^{-10}$, $\Omega_{1}=1$, $\Omega_{2}=1.1$, $\Omega_{c}=1$, $\Delta_{1}=\Delta_{2}=0.001$, $\Delta_{c}=-0.2$, $\gamma_{1}=\gamma_{2}=\gamma_{c}=0.1$ and $\phi=0.1320$ radian.
@@ II.1 Solution in mean-field approximation
Due to the analytical complexity of solving the nonlinear Eqs. (7) and (8), we employ the mean-field approximation to linearize these equations. In the limit of large excitation of the bosonic modes, the relevant operators can be expressed as the sum of their mean values and quantum fluctuations near the mean values, i.e., ${m}_{i}\rightarrow\left\langle{m_{i}}\right\rangle+\delta{m}_{i}={\alpha}_{i}+\delta{m}_{i}$ and ${a}\rightarrow\left\langle{a}\right\rangle+\delta{a}={\beta}+\delta{a}$. The equations for these mean values $\alpha_{i}$ and $\beta$ are written as,
@@
$\quad-g_{i}|\alpha_{i}|^{2}\beta A_{i}e^{-\iota t(B_{i}-C_{i})}+\Omega_{i}A_{i}e^{-\iota t(D_{i}+C_{i})}\alpha_{i}^{2}\bigg{]}$
$\dot{\beta}$ $=-\frac{\gamma_{c}}{2}\beta-\iota\left(g_{1}e^{\iota t(B_{1}-C_{1})}\alpha_{1}+g_{2}e^{\iota t(B_{2}-C_{2})}\alpha_{2}\right)-\iota\Omega_{c}e^{\iota t\Delta_{c}}$ (12)
In the limit of negligible higher order fluctuations, we obtain the linearized equations for the quantum fluctuations, as well, given by
$\delta\dot{m}_{i}$ $=$ $P_{i}\delta{m}_{i}+Q_{i}\delta{m}_{i}^{\dagger}+R_{i}\delta{a}+S_{i}\delta{a}^{\dagger}+\sqrt{\gamma_{i}}m_{\text{in}}^{(i)}+F_{i}\;,$
$\delta\dot{a}$ $=$ $\sum_{i=1}^{2}\left(U_{i}\delta{m}_{i}+W_{i}\delta{m}_{i}^{\dagger}\right)+T\delta{a}+\sqrt{\gamma_{c}}a_{\text{in}}+F_{3}\;,$ (13)
where $\delta m_{i}=\frac{\delta q_{i}+\iota\,\delta p_{i}}{\sqrt{2}},\quad i\in 1,2$  and $\delta a=\frac{\delta x+\iota\,\delta y}{\sqrt{2}}$ The expressions of the parameters $P_{i}$, $Q_{i}$, $R_{i}$, $S_{i}$, $U_{i}$, $W_{i}$, $T$, $F_{i}$, and $F_{3}$ listed in Appendix A.
@@
To ascertain the degree of quantum synchronization exhibited between oscillators, we adopt a figure of merit originally proposed by Mari et al. PhysRevLett.111.103605 . In the case, when the quantum fluctuations can be ignored, the classical synchronization prevails and may be quantified as
@@
In quantum regime, when quantum fluctuations cannot be ignored, the above variables need to be redefined with respect to their mean values as follows:
@@
which lies between 0 and 1. Here $S_{q}^{\phi}(t)=1$ indicates perfect phase-dependent synchronization and a null $S_{q}^{\phi}(t)$ denotes no synchronization.

### v2

@@ I INTRODUCTION
While extensively studied in classical contexts, recent research has extended this field into the quantum domain buvca2022algebraic; e26050415. Synchronization has been observed or predicted in a variety of quantum systems, ranging from subatomic particles to mechanical resonators manzano2013avoiding; manzano2013synchronization and electrodynamic setups buvca2022algebraic; eneriz2019degree. These systems often involve multiple interacting modes that exhibit nonlinear dynamics and are subject to strong external drives. For instance, recent experiments demonstrated quantum synchronization of a single trapped ion qubit with an external signal zhang2023quantum. Quantum synchronization has also been investigated across various other platforms, namely, van der Pol oscillators giorgi2012quantum, Josephson junction arrays wiesenfeld1996synchronization, spin torque nano-oscillators kaka2005mutual, and optomechanical systems holmes2012synchronization; zhang2012synchronization; heinrich2011collective; PhysRevLett.111.103605; shim2007synchronized. It has been shown that quantum fluctuations, nonclassical correlations, and phase transitions play critical roles in quantum synchronization phenomena buvca2022algebraic; eneriz2019degree; PhysRevA.102.032414. Recently, we have developed a theoretical framework to analyze quantum synchronization between two finite-size spin chains that interact indirectly through a central spin chain ghildiyal2025quantum. This configuration results in effective nonlinear coupling, with dynamics influenced by both linear and nonlinear dissipation.
@@
In this paper, we focus on quantum synchronization between two magnon modes. Magnons are quanta of spin-wave excitations arising from the collective precession of localized spins princep2017full; PhysRevResearch.2.022027; ghasemian2023dissipative; kusminskiy2021cavity. It has recently been shown that propagating spin waves can mediate phase-tunable synchronization between spin Hall nano-oscillators kumar2025spin.
@@
The ability to couple magnons with cavity modes and qubits has further led to a plethora of work related to quantum information processing in cavity-magnonic systems YUAN20221. This development has made it possible to explore quantum synchronization on such a platform. For example, it is shown that magnon squeezing can significantly enhance entanglement between two indirectly coupled magnon modes in a magnomechanical system harraf2025quantum. Yang and his coworkers demonstrated asymmetric quantum synchronization between two magnon modes in a two-sublattice antiferromagnet coupled to a cavity yang2022asymmetric, highlighting a strong dependence on cavity resonance and input current direction. We also note that the magnetostriction in a cavity magnomechanical system enables robust synchronization between two mechanical modes, via nonlinear dispersive coupling PhysRevResearch.5.043197. Nonreciprocal quantum synchronization in antiferromagnet-cavity systems has been reported in yang2021implementationenhancementnonreciprocalquantum. On the contrary, in our work, we show that magnon-magnon synchronization is indeed possible, just via their common coupling with a single-mode cavity, without requiring any squeezing or magnetostriction. A strong driving field in the presence of an intermode Kerr nonlinear interaction can lead to complete quantum synchronization. We have established our results using detailed time-dependent dynamics and the steady state behaviour of the magnon modes.
@@
The Yttrium Iron Garnet (YIG) spheres host well-defined magnon modes governed by magnetostatic solutions to the Landau-Lifshitz-Gilbert equation, thanks to their spherical structure. These spheres are often preferred in the studies of magnons due to their low magnetic damping and high coherence, enabling long-lived spin excitations. By placing the YIG spheres inside a high-quality microwave cavity, strong coupling between the magnon modes and the cavity photons can be achieved. This interaction leads to the formation of hybridized states known as cavity magnon polaritons, which are important for quantum technologies involving signal processing, coherent transduction, and nonlinear dynamics zhang2017observation. In this paper, we focus on quantum synchronization between two magnon modes of two such spatially separated YIG spheres, placed inside a cavity. The cavity-mediated mechanism enables the emergence of quantum synchronization in the presence of dissipation and detuning. We analyze the effects of key parameters, such as coupling strength, detuning, and nonlinear interactions between magnons, and identify the conditions under which stable synchronization emerges. Unlike other relevant works tyagi2024noise; bittner2025noise, we have employed the input-output formalism, instead of the master equation technique, to analyze the effect of decoherence.
@@ II MODEL AND EQUATION OF MOTION
We consider two macroscopic YIG spheres, trapped inside a single-mode microwave cavity. Each sphere hosts many magnon modes. The lowest-order mode, known as the Kittel mode, corresponds to the collective spin ground state, featuring uniform spin precession across the sample princep2017full. The higher-order modes exhibit non-uniform spatial distributions with specific angular and radial profiles. Here, we consider the YIG spheres in their respective Kittel mode. The cavity mode with the annihilation operator $a$ interacts coherently with these magnon modes with respective annihilation operators $m_{1}$ and $m_{2}$. This setup allows us to create an indirect coupling between the magnon modes through the cavity.
Figure 1: Schematic representation of two YIG (Yttrium Iron Garnet) spheres, ${m}_{1}$ and ${m}_{2}$, coupled via a common microwave cavity mode, a{a}.
The cavity is driven by an electromagnetic field of frequency $\omega_{L}$. Typically, the magnetic component of this driving field couples with the magnon mode of the macroscopic ensemble inside the sphere, via Zeeman interaction. The interaction Hamiltonian between the magnon modes and the cavity can then be written as PhysRevB.94.224410; PhysRevLett.104.077202
@@
Here $\omega_{c}$ is the cavity resonance frequency and $\omega_{m_{i}}$ is the frequency of the $i$th magnon mode. The $\Omega_{c}$ and $\Omega_{i}$ denote the Rabi frequencies, describing the amplitudes of the driving fields with respective frequencies $\omega_{L}$ and $\omega_{d}^{(i)}$, respectively. The term containing the Kerr coefficient $K_{i}$ arises from the magnetocrystalline anisotropy in YIG, which introduces intensity-dependent frequency shifts proportional to the magnon population $\langle m_{i}^{\dagger}m_{i}\rangle$ PhysRevB.94.224410; ganthya2022bistability. This term leads to a nonlinear phase-inducing evolution PhysRevB.94.224410; PhysRevResearch.1.023021 of the magnons. The Hamiltonian (2) contains the terms for cavity self-energy and the cavity driving, while Eq. (3) is a sum of magnon self-energy, nonlinear energy shift, and magnon driving term, respectively. In the rotating frame with respect to the driving frequencies $\omega_{L}$ and $\omega_{d}^{(i)}$, the explicit time-dependence in the Hamiltonian gets eliminated, leading to the following form of the total Hamiltonian:
@@
where $\Delta_{c}=\omega_{c}-\omega_{L}$ and $\Delta_{i}=\omega_{m_{i}}-\omega_{d}^{(i)}$ are the detunings of the cavity mode and the $i$th magnon mode, respectively, from the corresponding driving field.
Incorporating the interaction picture with respect to the energy terms containing $a^{\dagger}a$ and $m_{i}^{\dagger}m_{i}$ and using the Baker–Campbell–Hausdorff formalism, the modified Hamiltonian $H^{\prime}(t)$ can be expressed as $(\iota=\sqrt{-1})$
@@
$\quad-g_{i}m_{i}^{\dagger}A_{i}e^{-\iota t(B_{i}-C_{i})}m_{i}a+\Omega_{i}A_{i}e^{-\iota t(D_{i}+C_{i})}m_{i}^{2}\bigg],\$
@@
where $A_{i}=2\iota K_{i}t$ ($i\in 1,2$) is linearly proportional to the time $t$.
@@
Figure 2: Limit-cycle trajectories in the $q_{1}\leftrightharpoons p_{1}$ (red) and $q_{2}\leftrightharpoons p_{2}$ (blue) spaces (a, d, g), and variation of the mean values $\bar{q}_{1}$ (red) and $\bar{q}_{2}$ (blue) (b, e, h), along with mean values $\bar{p}_{1}$ (red) and $\bar{p}_{2}$ (blue) (c, f, i). The chosen parameters are: (a–c) $\Omega_{1}=1$, $\Omega_{2}=1.00001$, $g_{1}=g_{2}=0.1$. (d–f) $\Omega_{1}=1$, $\Omega_{2}=1.1$, $g_{1}=g_{2}=0.1$. (g–i) $\Omega_{1}=1$, $\Omega_{2}=1.00001$, $g_{1}=0.2$, $g_{2}=0.1$. All other parameters are identical across cases: $\Omega_{c}=1$, $\Delta_{1}=\Delta_{2}=0.001$, $\Delta_{c}=-0.2$, $K_{1}=K_{2}=10^{-10}$, and $\gamma_{1}=\gamma_{2}=\gamma_{c}=0.1$. All parameters are normalized with respect to $\Omega_{1}$ with the initial conditions $(\bar{q}_{1},\bar{p}_{1})=(1,0)=(\bar{q}_{2},\bar{p}_{2})$. The limit cycle trajectories are shown till $\Omega_{1}t=10^{5}$.
Figure 3: Limit-cycle trajectories in the $q_{1}\leftrightharpoons p_{1}$ (red) and $q_{2}\leftrightharpoons p_{2}$ (blue) spaces (a, d, g); variation of the mean values $\bar{q}_{1}$ (red), $\bar{q}_{2}$ (blue) (b, e, h); and $\bar{p}_{1}$ (red), $\bar{p}_{2}$ (blue) (c, f, i). The parameters for each case are: (a–c) $\Omega_{1}=1$, $\Omega_{2}=1.00001$$\Delta_{1}=\Delta_{2}=0.001$. (d–f) $\Omega_{1}=1$, $\Omega_{2}=1.1$, $\Delta_{1}=\Delta_{2}=0.001$. (g–i) $\Omega_{1}=1$, $\Omega_{2}=1.00001$, $\Delta_{1}=0.001$, $\Delta_{2}=0.0025$. Other parameters (common to all cases): $\Omega_{c}=1$, $\Delta_{c}=-0.2$, $g_{1}=g_{2}=0.5$, $K_{1}=K_{2}=10^{-10}$, and $\gamma_{1}=\gamma_{2}=\gamma_{c}=0.1$. All parameters are normalized with respect to $\Omega_{1}$ with the initial conditions $(\bar{q}_{1},\bar{p}_{1})=(1,0)$ and $(\bar{q}_{2},\bar{p}_{2})=(2,0)$. The limit cycle trajectories are shown till $\Omega_{1}t=10^{5}$.
Figure 4: (a) Variation of $S_{q}^{\phi}$, with respect to time $t$ , (b) variation of $\bar{S}_{q}^{\phi}$, with respect to the mean phonon number $\bar{n}_{m}$ of the environment. The other parameters are $g_{1}=g_{2}=0.1$, $K_{1}=K_{2}=10^{-10}$, $\Omega_{1}=1$, $\Omega_{2}=1.1$, $\Omega_{c}=1$, $\Delta_{1}=\Delta_{2}=0.001$, $\Delta_{c}=-0.2$, $\gamma_{1}=\gamma_{2}=\gamma_{c}=0.1$, and $\phi=0.1320$ radian. We have chosen an initial condition $(\bar{q}_{1},\bar{p}_{1})=(1,1)=(\bar{q}_{2},\bar{p}_{2})$.
@@ II.1 Solution in mean-field approximation
Due to the analytical complexity of solving the nonlinear Eqs. (7) and (8), we employ the mean-field approximation to linearize these equations. In the limit of large excitation of the bosonic modes, the relevant operators can be expressed as the sum of their mean values and quantum fluctuations near the mean values, i.e., ${m}_{i}\rightarrow\bar{m}_{i}+\delta{m}_{i}={\alpha}_{i}+\delta{m}_{i}$ and ${a}\rightarrow\bar{a}+\delta{a}={\beta}+\delta{a}$. The equations for these mean values $\alpha_{i}$ and $\beta$ are written as,
@@
$\quad-g_{i}|\alpha_{i}|^{2}\beta A_{i}e^{-\iota t(B_{i}-C_{i})}+\Omega_{i}A_{i}e^{-\iota t(D_{i}+C_{i})}\alpha_{i}^{2}\bigg]\;,$
$\dot{\beta}$ $=-\frac{\gamma_{c}}{2}\beta-\iota\left(g_{1}e^{\iota t(B_{1}-C_{1})}\alpha_{1}+g_{2}e^{\iota t(B_{2}-C_{2})}\alpha_{2}\right)-\iota\Omega_{c}e^{\iota t\Delta_{c}}\;.$ (12)
In the limit of negligible higher-order fluctuations, we obtain the linearized equations for the quantum fluctuations, as well, given by
$\delta\dot{m}_{i}$ $=$ $P_{i}\delta{m}_{i}+Q_{i}\delta{m}_{i}^{\dagger}+R_{i}\delta{a}+S_{i}\delta{a}^{\dagger}+\sqrt{\gamma_{i}}m_{\text{in}}^{(i)}\;,$
$\delta\dot{a}$ $=$ $\sum_{i=1}^{2}\left(U_{i}\delta{m}_{i}+W_{i}\delta{m}_{i}^{\dagger}\right)+T\delta{a}+\sqrt{\gamma_{c}}a_{\text{in}}\;,$ (13)
where $\delta m_{i}=\frac{\delta q_{i}+\iota\,\delta p_{i}}{\sqrt{2}},\quad i\in 1,2$  and $\delta a=\frac{\delta x+\iota\,\delta y}{\sqrt{2}}$ . The expressions of the parameters $P_{i}$, $Q_{i}$, $R_{i}$, $S_{i}$, $U_{i}$, $W_{i}$, $T$, $F_{i}$, and $F_{3}$ are listed in Appendix A.
@@
To ascertain the degree of quantum synchronization between oscillators, we adopt a figure of merit originally proposed by Mari et al. PhysRevLett.111.103605. In the case when the quantum fluctuations can be ignored, the classical synchronization prevails and may be quantified as
@@
In the quantum regime, when quantum fluctuations cannot be ignored, the above variables need to be redefined with respect to their mean values as follows:
@@
which lies between 0 and 1. Here, $S_{q}^{\phi}(t)=1$ indicates phase-dependent complete synchronization, and a null $S_{q}^{\phi}(t)$ denotes no synchronization.